Given that μXμX = 2 and σ2X=σX2=15,

a) Let Y = 6X + 143 and find σ2YσY2 : , and μYμY :

b) Let Z = 6X + 57 and find σ2ZσZ2:, and μZμZ:
Answer #1

(a)

Given:

mu _{X}=2 (1)

= 15 (2)

Y = 6X + 143 (3)

By Theorem:

(i)

E(aX + b) = aE(X) + b (4)

(ii)

Var(aX + b) = a-Var(X) (5)

(iii)

E(aX+bY)=aE(X)+bE(Y) (6)

(iv)

Var(aX + bY) = a2 Var(X) + b2Var(Y) (7)

From (3), we get:

a = 6

b = 143

Substituting in (5), we get:

sigma^{2} _{Y}=6^{2} imes Var(X) (8)

Substituting (2), equation(8) becomes:

sigma^{2} _{Y}=6^{2} imes 15=540

Substituting in (4), we get:

E(Y) 6E(X) 143 (9)

Substituting (1), equation (9) becomes:

A. (6 × 2) 143-155 (10)

(b)

Given:

mu _{X}=2 (11)

= 15 (12)

Y = 6X + 57 (13)

By Theorem:

(i)

E(aX + b) = aE(X) + b (14)

(ii)

Var(aX + b) = a-Var(X) (15)

(iii)

E(aX+bY)=aE(X)+bE(Y) (16)

(iv)

Var(aX + bY) = a2 Var(X) + b2Var(Y) (17)   

From (13), we get:

a = 6

b = 57

Substituting in (15), we get:

sigma^{2} _{Y}=6^{2} imes Var(X) (18)

Substituting (2), equation(8) becomes:

sigma^{2} _{Y}=6^{2} imes 15=540

Substituting in (4), we get:

E(Y)=6E(X)+57 (19)

Substituting (11), equation (19) becomes:

mu _{Y}=(6 imes 2)+57=69 (20)

Thus, Answers to Questions asked:

(a)

sigma^{2} _{Y}=540

mu _{Y}=155

(b)

sigma^{2} _{Y}=540

mu _{Y}=69